# FastDINOv2: Frequency Based Curriculum Learning Improves Robustness and Training Speed

### v1

@@ Abstract
Large-scale vision foundation models such as DINOv2 boast impressive performances by leveraging massive architectures and training datasets. But numerous scenarios require practitioners to reproduce those pre-training solutions, such as on private data, new modalities, or simply for scientific questioning–which is currently extremely demanding computation-wise. We thus propose a novel pre-training strategy for DINOv2 that simultaneously accelerates convergence–and strengthens robustness to common corruptions as a by-product. Our approach involves a frequency filtering curriculum–low-frequency being seen first–and the Gaussian noise patching augmentation. Applied to a ViT-B/16 backbone trained on ImageNet-1K, while pre-training time and FLOPs are reduced by 1.6×1.6$\times$1.6 × and 2.25×2.25$\times$2.25 ×, our method still achieves matching robustness in corruption benchmarks (ImageNet-C) and maintains competitive linear probing performance compared with baseline. This dual benefit of efficiency and robustness makes large-scale self-supervised foundation modeling more attainable, while opening the door to novel exploration around data curriculum and augmentation as means to improve self-supervised learning models robustness. The code is available at https://github.com/KevinZ0217/fast_dinov2
@@
Training Method Training Time (days)()($\downarrow$)( ↓ ) ImageNet-1K()($\uparrow$)( ↑ ) ImageNet-C ()($\downarrow$)( ↓ ) GFLOPs ()($\downarrow$)( ↓ )
DINOv2 16.64(l40s) 77.8% 56.5% 493.76
FastDINOv2 10.32(l40s) 76.2% 56.7% 219.92
@@ 1 Introduction
Models such as DINOv2 [22] and CLIP [25], built on Vision Transformer (ViT) backbones [11], achieve remarkable performance, generalization, and even upstream robustness  [39] [24]. These advances stem largely from self-supervised learning (SSL)—a paradigm in which models learn useful representations from unlabeled data by solving proxy tasks (e.g., contrastive matching, predicting masked patches) rather than relying on costly manual annotation [12]. SSL’s appeal lies in its ability to leverage the massive, ever-growing pools of raw images available online; by discovering structure in data itself, SSL yields features that transfer effectively to downstream tasks such as classification, objection detection, and segmentation, often matching or exceeding their fully supervised counterparts [9] [13].
However, reproducing these recent breakthroughs typically demands huge compute: large ViT variants (e.g., ViT-G) trained for hundreds of epochs on billions of images [4], multi-GPU clusters, and sophisticated optimization recipes. This resource barrier can put state-of-the-art SSL out of reach for many academic labs and startups, limiting both reproducibility and further innovation. Moreover, while pre-trained SSL models frequently demonstrate excellent performance on clean benchmarks, robustness to real-world distribution shifts—common corruptions, sensor noise, weather effects—remains crucial for safety-critical applications such as medical imaging [2] [30]or autonomous driving [4]. Although some studies have shown that SSL can improve robustness compared to supervised pre-training [18] [3] [21], most SSL methods do not explicitly optimize for robustness, and existing robust pre-training often compounds the compute demands [26] [35].
In particular, recent large-scale SSL models exhibit an emergent robustness only when trained at extreme scales: for instance, DINOv2 and related methods require model sizes upward of 86 millions of parameters for ViT-B and datasets of similar magnitude, such as LVD-142M and LAION-5B [28], to yield strong resistance to corruptions. This scale-driven robustness is appealing in principle, but is prohibitively expensive in practice for researchers without access to massive compute. Consequently, there is a pressing need to design compute-efficient SSL training method that still deliver robustness guarantees.
@@
  • The proposed curriculum accelerates convergence on ImageNet-1K [27] with DINOv2 and a ViT-B backbone, reducing pretraining time by 1.66×1.$6$ and FLOPs by 2.25×2.25$\times$2.25 ×, while maintaining matching robustness and competitive clean linear probing accuracy.
@@ 2 Related Work
By ordering the example from the easiest to the hardest, curriculum learning yields a monotonic increase in the rate of convergence under stochastic gradient descent [34]. It has also been proven that an ideal curriculum can effectively smooth the optimization landscape without changing the location of global minima [14]. Thus, defining the level of difficulty effectively is essential for a powerful curriculum. Gradient-based training of deep networks systematically fits the low-frequency components of a target function first, and gradually capture high-frequency features as training proceeds [36]. For vision transformer, specifically, the multi-head self-attention acts as a low-pass filter, attenuating high-frequency signals while preserving low-frequency components [23].
The coarse-to-fine paradigm in computer vision motivates training models first in images with reduced information, then on their full-resolution counterparts. Therefore, it is natural to define low-frequency components from an image as easy examples, while full image with both low- and high-frequency features as more difficult ones. From a Fourier perspective, natural image content is concentrated in the low-frequency domain [37], so downsampling preserves most of the information while reducing computational cost. Several works incorporate small images or low-frequency components into the training pipeline to accelerate ViT convergence: RECLIP[19] applies this curriculum to pre-train CLIP, and EfficientTrain++ [33] generalizes it for various ViT-based models, including MoCo [15] and MAE [16]. However, the impact of this approach on model robustness remains unclear. Moreover, as a general method, EfficientTrain++ evaluated it on DINO [7], but did not observe convergence speedup. In this paper, we revisit the curriculum learning applied to DINOv2. Besides the acceleration in training convergence, we also discover that this data curriculum unexpectedly biases model toward high-frequency information.
@@
Robustness research often adopts a frequency-domain perspective to examine how image corruptions correspond to the frequency spectrum [37]. In natural images, low-frequency components dominate, carrying most structural information, so high-frequency corruptions, such as Gaussian or shot noise, primarily disrupt fine details like edges, while low-frequency corruptions, including contrast shifts, brightness changes, or fog, modify broader patterns. Recent works have studied the need to model the noise as part of the data augmentation strategies–justifying the use of such corruptions as means to improve downstream robustness [31]. Different data augmentations introduce frequency biases that enhance robustness to specific corruptions but reveal distinct strengths and weaknesses [8]. For example, Gaussian noise augmentation strengthens robustness to high-frequency perturbations, producing a low-frequency-biased model, but often degrades performance on medium-frequency distortions like motion blur or Gaussian blur [37]. Similarly, CutOut [10] augmentation, which masks image regions to emphasize high-frequency cues such as edges or textures, encourages models to rely on features that are easily corrupted by noise or blur, limiting robustness gains. Adversarial training, on the other hand, biases models toward low-frequency features, enhancing resilience to certain perturbations but compromising performance on low-frequency corruptions [38] [6]. While these augmentations offer specific strengths, they often impair generalization by prioritizing certain frequency bands over others. In this work, we address this tradeoff by integrating Gaussian noise patching [20], which applies noise to localized image patches to balance robustness to high-frequency corruptions with preserved accuracy on clean images. This approach complements the high-frequency bias induced by our data curriculum, fostering a more balanced model across the frequency spectrum, though careful evaluation is needed to mitigate potential weaknesses on medium-frequency corruptions.
@@ 3.1 Low-Frequency Feature Extraction
Low-frequency features can be extracted through multiple approaches. While methods like EfficientTrain++ [33] employ Fourier transform with high-frequency filtering for this purpose, we opt for a simpler and more computationally efficient strategy. In our pipeline, we use downsampling as a lightweight proxy for low-frequency extraction, applied after DINOv2’s standard image cropping preprocessing. In DINOv2’s teacher-student framework, the teacher network processes global crops, which preserve low-frequency structural information, while the student network handles local crops, which retain high-frequency details. The cropping operation is defined as follows:
@@ 3.2 Frequency Based Curriculum Learning
Curriculum learning [5] improves model performance by progressively training on simpler samples before advancing to more complex ones. In our approach, we define low-frequency components of images—representing coarse, large-scale patterns—as the easier samples, gradually progressing to harder examples represented by the original images. Our curriculum consists of two stages. For the first 75% of training epochs, the model is trained exclusively on these low-frequency components. Subsequently, we apply a restarting mechanism by resetting the Adam optimizer’s training dynamics. In the second stage, we train DINOv2 on original images containing both low- and high-frequency information for the remaining epochs. To ensure training stability, we maintain a fixed batch size across both stages.
@@ 4.1 Datasets and Training Setup
For faster experimentation, we primarily use ImageNet-100 [32], a subset of ImageNet-1K [27]. The training set consists of 100 randomly selected classes from ImageNet-1K, with the first 500 images from each class. Similarly, the validation set contains the corresponding 100 classes from the original validation set, with 50 images per class. This results in a total of 50,000 training images and 5,000 validation images. For robustness evaluation, we use ImageNet-100-C, derived from ImageNet-C [17], which benchmarks model resilience to common corruptions. We maintain the exact image selection from the ImageNet-100 validation set across all corruption levels and types. Additionally, we employ ADE20K for semantic segmentation tasks. Finally, we scale our approach to full ImageNet-1K and evaluate robustness on ImageNet-C.
@@
For ImageNet-100 experiments, we use ViT-S/16 as the DINOv2 backbone with a total batch size of 40, distributed across 4 GPUs (10 per GPU). Experiments for large batch size in first stage of training are in Section A.1. We adopt positional embedding with interpolation for adapting to image resolution shift between two training stages. The results of applying positional embedding with or without interpolation across different embedding sizes are in Section A.2. We train baseline models for 500 epochs and training curriculum experiments for 200 epochs, ensuring the baseline converges to optimal performance. All ImageNet-100 experiments use a fixed epoch length of 1,250 iterations. For ImageNet-1K experiments, we employ ViT-B/16 with a total batch size of 512 (128 per GPU), with epoch length of 2,500 iterations. The baseline and FastDINOv2 are trained for 250 and 200 epochs on ImageNet-1K, respectively. Following the official DINOv2 implementation, we use AdamW optimizer with square root learning rate scaling based on batch size, yielding a base learning rate of $7.9\times 10^{-4}$. For linear probing evaluation, we use a batch size of 128 with 12.5k total iterations. All training runs are distributed across 4 L40S GPUs, while evaluations use either A6000 or A5500 GPUs.
@@ 4.2 Linear Probing Reveals Fast Convergence Without Compromising Accuracy
Notably, the 112-224 curriculum achieves the baseline’s 250-epoch accuracy by epoch 200, demonstrating 20% faster convergence in training epochs and 1.44×1.$4$ acceleration in training time. We additionally validate the framework’s flexibility by testing alternative first-stage resolutions in Table 2, observing consistent efficiency gains across configurations.
@@
Table 2: Linear probing accuracy on ImageNet-100. Our data curriculum pipeline matches DINOv2 performance while achieving a 1.44$\times$ convergence speed-up. 112-224 means that the FastDINOv2 training utilizes the 112-224 data curriculum; same applies to the remaining training methods.
Training method Accuracy Training epoch Training time (l40s hours)
@@ 4.3 Data Curriculum Induces High-Frequency Feature Bias
To better interpret this shift in robustness, we group corruption types in Table 3 based on their dominant frequency characteristics, following a Fourier analysis approach [37]:
@@ 4.4 Spectral Balance: Combining Curriculum and Noise Patching
We identify Gaussian noise augmentation [1] as a method to introduce low-frequency bias, demonstrated by enhanced robustness against high-frequency noise corruptions. However, its direct application decreases clean accuracy due to excessive low-frequency emphasis. Gaussian noise patching mitigates this by selectively applying localized noise injection alongside preserved clean regions, thereby maintaining discriminative feature learning.
@@ 4.7 Frequency Based Analysis and Grad-CAM Visual Explanation
In addition to the evaluations using ImageNet-100-C and linear probing, we explore whether different training paradigms influence the error sensitivity of models across all frequency bands—not only those in ImageNet-100-C—and the regions the model prioritizes during predictions. To this end, we generate Grad-CAM visualizations [29] and Fourier spectral heatmaps following the methods in [37].
@@
Figure 3: Grad-CAM maps for DINOv2 baseline and FastDINOv2. The first row images are from class "house finch, linnet, Carpodacus mexicanus", and the second row from "goose". With the data curriculum, model can better capture the contour of the object. See more examples in  A.4
@@ 4.8 Extending to ViT-B and ImageNet-1K: Balancing Efficiency and Robustness
Table 1 reports the evaluation results and training costs for the baseline and our method. Although training time is reduced by 1.6$\times$ and FLOPs by 2.25$\times$, our method achieves competitive linear probing accuracy. Moreover, testing on ImageNet-C with a frozen backbone and linear classifier shows a comparable error rate despite a slight drop in clean accuracy, demonstrating the effectiveness of our method in enhancing robustness.
@@
Table 7: Comparison of training epochs and max memory consumptions on a single l40s GPU with 48GB memory. Both DINOv2 and FastDINOv2 apply ViT-B backbone trained on ImageNet-1K.

### v2

@@ Abstract
Large-scale vision foundation models such as DINOv2 boast impressive performances by leveraging massive architectures and training datasets. But numerous scenarios require practitioners to reproduce those pre-training solutions, such as on private data, new modalities, or simply for scientific questioning–which is currently extremely demanding computation-wise. We thus propose a novel pre-training strategy for DINOv2 that simultaneously accelerates convergence–and strengthens robustness to common corruptions as a by-product. Our approach involves a frequency filtering curriculum–low-frequency being seen first–and the Gaussian noise patching augmentation. Applied to a ViT-B/16 backbone trained on ImageNet-1K, while pre-training time and FLOPs are reduced by $1.6\times$ and $2.25\times$, our method still achieves matching robustness in corruption benchmarks (ImageNet-C) and maintains competitive linear probing performance compared with baseline. This dual benefit of efficiency and robustness makes large-scale self-supervised foundation modeling more attainable, while opening the door to novel exploration around data curriculum and augmentation as means to improve self-supervised learning models robustness. The code is available at https://github.com/KevinZ0217/fast_dinov2
@@
Training Method Training Time (days)()($\downarrow$) ImageNet-1K()($\uparrow$) ImageNet-C $(\downarrow)$ GFLOPs $(\downarrow)$
DINOv2 16.64 (NVIDIA L40S) 77.8% 56.5% 493.76
FastDINOv2 10.32 (NVIDIA L40S) 76.2% 56.7% 219.92
@@ 1 Introduction
Models such as DINOv2 [dinov2] and CLIP [radford2021clip], built on Vision Transformer (ViT) backbones [dosovitskiy2021image], achieve remarkable performance, generalization, and even upstream robustness  [zhou2022robustnessvit] [paul2022vision]. These advances stem largely from self-supervised learning (SSL)—a paradigm in which models learn useful representations from unlabeled data by solving proxy tasks (e.g., contrastive matching, predicting masked patches) rather than relying on costly manual annotation [ericsson2022selfsupervised]. SSL’s appeal lies in its ability to leverage the massive, ever-growing pools of raw images available online; by discovering structure in data itself, SSL yields features that transfer effectively to downstream tasks such as classification, objection detection, and segmentation, often matching or exceeding their fully supervised counterparts [chen2020simclr] [grill2020byol].
However, reproducing these recent breakthroughs typically demands huge compute: large ViT variants (e.g., ViT-G) trained for hundreds of epochs on billions of images [barin2024beview], multi-GPU clusters, and sophisticated optimization recipes. This resource barrier can put state-of-the-art SSL out of reach for many academic labs and startups, limiting both reproducibility and further innovation. Moreover, while pre-trained SSL models frequently demonstrate excellent performance on clean benchmarks, robustness to real-world distribution shifts—common corruptions, sensor noise, weather effects—remains crucial for safety-critical applications such as medical imaging [baharoon2023towards] [song2024dino_reg]or autonomous driving [barin2024beview]. Although some studies have shown that SSL can improve robustness compared to supervised pre-training [hendrycks2019using] [balestriero2023cookbook] [lu2025ditch], most SSL methods do not explicitly optimize for robustness, and existing robust pre-training often compounds the compute demands [reed2022self] [xu2020adversarial].
In particular, recent large-scale SSL models exhibit an emergent robustness only when trained at extreme scales: for instance, DINOv2 and related methods require model sizes upward of 86 millions of parameters for ViT-B and datasets of similar magnitude, such as LVD-142M and LAION-5B [schuhmann2022laion5b], to yield strong resistance to corruptions. This scale-driven robustness is appealing in principle, but is prohibitively expensive in practice for researchers without access to massive compute. Consequently, there is a pressing need to design compute-efficient SSL training method that still deliver robustness guarantees.
@@
  • The proposed curriculum accelerates convergence on ImageNet-1K [russakovsky2015imagenet] with DINOv2 and a ViT-B backbone, reducing pretraining time by $1.66\times$ and FLOPs by $2.25\times$, while maintaining matching robustness and competitive clean linear probing accuracy.
@@ 2 Related Work
By ordering the example from the easiest to the hardest, curriculum learning yields a monotonic increase in the rate of convergence under stochastic gradient descent [weinshall2018curriculum]. It has also been proven that an ideal curriculum can effectively smooth the optimization landscape without changing the location of global minima [hacohen2019power]. Thus, defining the level of difficulty effectively is essential for a powerful curriculum. Gradient-based training of deep networks systematically fits the low-frequency components of a target function first, and gradually capture high-frequency features as training proceeds [xu2019frequency]. For vision transformer, specifically, the multi-head self-attention acts as a low-pass filter, attenuating high-frequency signals while preserving low-frequency components [park2022how].
The coarse-to-fine paradigm in computer vision motivates training models first in images with reduced information, then on their full-resolution counterparts. Therefore, it is natural to define low-frequency components from an image as easy examples, while full image with both low- and high-frequency features as more difficult ones. From a Fourier perspective, natural image content is concentrated in the low-frequency domain [yinetal2019fourier], so downsampling preserves most of the information while reducing computational cost. Several works incorporate small images or low-frequency components into the training pipeline to accelerate ViT convergence: RECLIP[li2023reclip] applies this curriculum to pre-train CLIP, and EfficientTrain++ [wang2024efficienttrainplusplus] generalizes it for various ViT-based models, including MoCo [he2020momentumcontrast] and MAE [he2022maskedautoencoders]. However, the impact of this approach on model robustness remains unclear. Moreover, as a general method, EfficientTrain++ evaluated it on DINO [caron2021emerging], but did not observe convergence speedup. In this paper, we revisit the curriculum learning applied to DINOv2. Besides the acceleration in training convergence, we also discover that this data curriculum unexpectedly biases model toward high-frequency information.
@@
Robustness research often adopts a frequency-domain perspective to examine how image corruptions correspond to the frequency spectrum [yinetal2019fourier]. In natural images, low-frequency components dominate, carrying most structural information, so high-frequency corruptions, such as Gaussian or shot noise, primarily disrupt fine details like edges, while low-frequency corruptions, including contrast shifts, brightness changes, or fog, modify broader patterns. Recent works have studied the need to model the noise as part of the data augmentation strategies–justifying the use of such corruptions as means to improve downstream robustness [van_assel2025joint_embedding]. Different data augmentations introduce frequency biases that enhance robustness to specific corruptions but reveal distinct strengths and weaknesses [chan2022frequencybias]. For example, Gaussian noise augmentation strengthens robustness to high-frequency perturbations, producing a low-frequency-biased model, but often degrades performance on medium-frequency distortions like motion blur or Gaussian blur [yinetal2019fourier]. Similarly, CutOut [devries2017cutout] augmentation, which masks image regions to emphasize high-frequency cues such as edges or textures, encourages models to rely on features that are easily corrupted by noise or blur, limiting robustness gains. Adversarial training, on the other hand, biases models toward low-frequency features, enhancing resilience to certain perturbations but compromising performance on low-frequency corruptions [zhang2024mitigating] [bu2023towards]. While these augmentations offer specific strengths, they often impair generalization by prioritizing certain frequency bands over others. In this work, we address this tradeoff by integrating Gaussian noise patching [lopes2019patchgaussian], which applies noise to localized image patches to balance robustness to high-frequency corruptions with preserved accuracy on clean images. This approach complements the high-frequency bias induced by our data curriculum, fostering a more balanced model across the frequency spectrum, though careful evaluation is needed to mitigate potential weaknesses on medium-frequency corruptions.
@@ 3.1 Low-Frequency Feature Extraction
Low-frequency features can be extracted through multiple approaches. While methods like EfficientTrain++ [wang2024efficienttrainplusplus] employ Fourier transform with high-frequency filtering for this purpose, we opt for a simpler and more computationally efficient strategy. In our pipeline, we use downsampling as a lightweight proxy for low-frequency extraction, applied after DINOv2’s standard image cropping preprocessing. In DINOv2’s teacher-student framework, the teacher network processes global crops, which preserve low-frequency structural information, while the student network handles local crops, which retain high-frequency details. The cropping operation is defined as follows:
@@ 3.2 Frequency Based Curriculum Learning
Curriculum learning [bengio2009curriculum] improves model performance by progressively training on simpler samples before advancing to more complex ones. In our approach, we define low-frequency components of images—representing coarse, large-scale patterns—as the easier samples, gradually progressing to harder examples represented by the original images. Our curriculum consists of two stages. For the first 75% of training epochs, the model is trained exclusively on these low-frequency components. Subsequently, we apply a restarting mechanism by resetting the Adam optimizer’s training dynamics. In the second stage, we train DINOv2 on original images containing both low- and high-frequency information for the remaining epochs. To ensure training stability, we maintain a fixed batch size across both stages.
@@ 4.1 Datasets and Training Setup
For faster experimentation, we primarily use ImageNet-100 [miniimagenet], a subset of ImageNet-1K [russakovsky2015imagenet]. The training set consists of 100 randomly selected classes from ImageNet-1K, with the first 500 images from each class. Similarly, the validation set contains the corresponding 100 classes from the original validation set, with 50 images per class. This results in a total of 50,000 training images and 5,000 validation images. For robustness evaluation, we use ImageNet-100-C, derived from ImageNet-C [hendrycks2019benchmarking], which benchmarks model resilience to common corruptions. We maintain the exact image selection from the ImageNet-100 validation set across all corruption levels and types. Additionally, we employ ADE20K for semantic segmentation tasks. Finally, we scale our approach to full ImageNet-1K and evaluate robustness on ImageNet-C.
@@
For ImageNet-100 experiments, we use ViT-S/16 as the DINOv2 backbone with a total batch size of 40, distributed across 4 GPUs (10 per GPU). Experiments for large batch size in first stage of training are in Section˜A.1. We adopt positional embedding with interpolation for adapting to image resolution shift between two training stages. The results of applying positional embedding with or without interpolation across different embedding sizes are in Section˜A.2. We train baseline models for 500 epochs and training curriculum experiments for 200 epochs, ensuring the baseline converges to optimal performance. All ImageNet-100 experiments use a fixed epoch length of 1,250 iterations. For ImageNet-1K experiments, we employ ViT-B/16 with a total batch size of 512 (128 per GPU), with epoch length of 2,500 iterations. The baseline and FastDINOv2 are trained for 250 and 200 epochs on ImageNet-1K, respectively. Following the official DINOv2 implementation, we use AdamW optimizer with square root learning rate scaling based on batch size, yielding a base learning rate of $7.9\times 10^{-4}$. For linear probing evaluation, we use a batch size of 128 with 12.5k total iterations. All training runs are distributed across 4 NVIDIA L40S GPUs, while evaluations use either NVIDIA A6000 or NVIDIA A5500 GPUs.
@@ 4.2 Linear Probing Reveals Fast Convergence Without Compromising Accuracy
Notably, the 112-224 curriculum achieves the baseline’s 250-epoch accuracy by epoch 200, demonstrating 20% faster convergence in training epochs and $1.44\times$ acceleration in training time. We additionally validate the framework’s flexibility by testing alternative first-stage resolutions in Table 2, observing consistent efficiency gains across configurations.
@@
[…]
Training method Accuracy Training epoch Training time
@@ 4.3 Data Curriculum Induces High-Frequency Feature Bias
To better interpret this shift in robustness, we group corruption types in Table 3 based on their dominant frequency characteristics, following a Fourier analysis approach [yinetal2019fourier]:
@@ 4.4 Spectral Balance: Combining Curriculum and Noise Patching
We identify Gaussian noise augmentation [akbiyik2023data] as a method to introduce low-frequency bias, demonstrated by enhanced robustness against high-frequency noise corruptions. However, its direct application decreases clean accuracy due to excessive low-frequency emphasis. Gaussian noise patching mitigates this by selectively applying localized noise injection alongside preserved clean regions, thereby maintaining discriminative feature learning.
@@ 4.7 Frequency Based Analysis and Grad-CAM Visual Explanation
In addition to the evaluations using ImageNet-100-C and linear probing, we explore whether different training paradigms influence the error sensitivity of models across all frequency bands—not only those in ImageNet-100-C—and the regions the model prioritizes during predictions. To this end, we generate Grad-CAM visualizations [selvaraju2017gradcam] and Fourier spectral heatmaps following the methods in [yinetal2019fourier].
@@
Figure 3: Grad-CAM maps for DINOv2 baseline and FastDINOv2. The first row images are from class "house finch, linnet, Carpodacus mexicanus", and the second row from "goose". With the data curriculum, model can better capture the contour of the object. See more examples in  A.7
@@ 4.8 Extending to ViT-B and ImageNet-1K: Balancing Efficiency and Robustness
[…]
@@
Table 7: Comparison of training epochs and max memory consumptions on a single NVIDIA L40S GPU with 48GB memory. Both DINOv2 and FastDINOv2 apply ViT-B backbone trained on ImageNet-1K.